# Solvent: liquidity verification of smart contracts

### v1

@@ 0.A.3 Use case: a 2-players lottery
We show below an implementation of the lottery protocol in Solvent; we then apply our tool to verify some relevant liveness properties. Intuitively, the protocol followed by honest players is the following:
@@
We implement the lottery protocol in LABEL:lst:lottery1 and LABEL:lst:lottery2. The expected liveness properties of the contract, formalised in LABEL:lst:lottery-properties, are the following:
@@
(app_tx_st.balance[player1] >= st.balance[player1] + 1)
@@
(app_tx_st.balance[player2] >= st.balance[player2] + 2)
@@
(app_tx_st.balance[xa] >= st.balance[xa] + st.balance)
@@
(app_tx_st.balance[player1] >= st.balance[player1] + 2)
@@
(app_tx_st.balance[player1] >= st.balance[player1] + 2) || (app_tx_st.balance[player2] >= st.balance[player2] + 2))

### v2

@@ 0.A.4 Use case: a 2-players lottery
We show below an implementation of the lottery protocol in Solvent; we then apply our tool to verify some relevant liquidity properties. Intuitively, the protocol followed by honest players is the following:
@@
We implement the lottery protocol in LABEL:lst:lottery1 and LABEL:lst:lottery2. The expected liquidity properties of the contract, formalised in LABEL:lst:lottery-properties, are the following:
@@
(<tx>player1.balance >= player1.balance + 1)
@@
(<tx>player2.balance >= player2.balance + 2)
@@
(<tx>xa.balance >= xa.balance + balance)
@@
(<tx>player1.balance >= player1.balance + 2)
@@
(<tx>player1.balance >= player1.balance + 2) ||
(<tx>player2.balance >= player2.balance + 2))